Equation in which M(x,y) and N(x,y) are Linear but not Homogeneous


Find the general solution of the following:

1.) (2x+3y-1)dx-4(x+1)dy=0

2.) (x-4y-3)dx-(x-6y-5)dy=0


1
Expert's answer
2021-06-10T05:29:30-0400

Let us find the general solution of the following:


1) "(2x+3y-1)dx-4(x+1)dy=0"


Let "x=u-1,\ y =v+1." Then "(2(u-1)+3(v+1)-1)du-4udv=0."

It follows that "(2u+3v)du-4udv=0."


Let "u=tv," then "du=tdv+vdt." Then "(2tv+3v)(tdv+vdt)-4tvdv=0."

It follows that "(2t+3)(tdv+vdt)-4tdv=0," and hence "(2t^2+3t)dv+(2t+3)vdt-4tdv=0" which is equivalent to "(2t^2-t)dv+(2t+3)vdt=0." Taking into account that "v=0", and thus "y=1" is not a solution, we have that "\frac{dv}{v}=-\frac{2t+3}{2t^2-t}dt=\frac{-2t-3}{(2t-1)t}dt=(\frac{3}{t}-\frac{8}{2t-1})dt," and thus "\int\frac{dv}{v}=\int(\frac{3}{t}-\frac{8}{2t-1})dt." It follows that "\ln v =3\ln t -4\ln|2t-1|+ln|C|," and hence "v=\frac{Ct^3}{(2t-1)^4}." Since "t=\frac{u}{v}=\frac{x+1}{y-1}," we have that "y-1=\frac{C(\frac{x+1}{y-1})^3}{(2(\frac{x+1}{y-1})-1)^4}=\frac{C(x+1)^3(y-1)}{(2x-y+3)^4}". Since "x=-1" is a solution, the general solution we can write in the form:


"(x+1)^3=C(2x-y+3)^4."



2) "(x-4y-3)dx-(x-6y-5)dy=0"


Let "x=u-1,\ y=v-1." Then "(u-1-4(v-1)-3)du-(u-1-6(v-1)-5)dv=0."

It follows that "(u-4v)du-(u-6v)dv=0."


Let "u=tv," then "du=tdv+vdt." Then "(tv-4v)(tdv+vdt)-(tv-6v)dv=0." Taking into account that "v=0", and thus "y=-1" is not a solution, we have that "(t-4)(tdv+vdt)-(t-6)dv=0." It follows that "(t^2-4t)dv+(t-4)vdt-(t-6)dv=0,"

and hence "(t^2-5t+6)dv+(t-4)vdt=0." We have that "\frac{dv}{v}=-\frac{t-4}{t^2-5t+6}dt=\frac{-t+4}{(t-2)(t-3)}dt=(\frac{1}{t-3}-\frac{2}{t-2})dt," and thus "\int\frac{dv}{v}=\int(\frac{1}{t-3}-\frac{2}{t-2})dt." Then "\ln |v|=\ln|t-3|-2\ln|t-2|+\ln|C|," and thus "v=\frac{C(t-3)}{(t-2)^2}." Taking into account that "t=\frac{u}{v}=\frac{x+1}{y+1}," we conclude that "y+1=\frac{C(\frac{x+1}{y+1}-3)}{(\frac{x+1}{y+1}-2)^2}=\frac{C(y+1)(x-3y-2)}{(x-2y-1)^2}." It follows that he general solution we can write in the form: "(x-2y-1)^2=C(x-3y-2)."